OpenStudy (anonymous):

Please Help! Find the domain and range. f(x)=log3(x-2)+3

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=\log_3(x-2)+3}\)

OpenStudy (solomonzelman):

you can't take a log of 0, or of any negative number. That would mean that x-2 has to be greater than 0

OpenStudy (solomonzelman):

For x-2 to be greater than 0 x-2>0 x>2

OpenStudy (solomonzelman):

got the domain part ?

OpenStudy (anonymous):

yes how would I get the range

OpenStudy (solomonzelman):

ok, for that follow along with me...

OpenStudy (solomonzelman):

Tell me what is (lets disregard the base 3, since it doesn't do anything in our case) what is \(\large\color{black}{ \displaystyle \log(0.1)}\) ?

OpenStudy (anonymous):

-1

OpenStudy (solomonzelman):

hint: \(\large\color{black}{ \displaystyle \log(0.1)}\) >> \(\large\color{black}{ \displaystyle \log(\frac{1}{10})}\) >> \(\large\color{black}{ \displaystyle \log(10^{-1})}\) >> \(\large\color{black}{ \displaystyle -\log(10)}\) >> \(\large\color{black}{ \displaystyle -1}\)

OpenStudy (solomonzelman):

correct

OpenStudy (solomonzelman):

What is \(\large\color{black}{ \displaystyle \log(0.01)}\) ?

OpenStudy (anonymous):

Its hard to read it seems you retyped a few things can you explain better

OpenStudy (solomonzelman):

I didn't re-type anything.

OpenStudy (solomonzelman):

I am just trying to get to the following: \(\large\color{black}{ \displaystyle \log(0.1)=\log(10^{-1})=-1}\) \(\large\color{black}{ \displaystyle \log(0.01)=\log(10^{-2})=-2}\) \(\large\color{black}{ \displaystyle \log(0.001)=\log(10^{-3})=-3}\) \(\large\color{black}{ \displaystyle \log(0.0001)=\log(10^{-4})=-4}\) \(\large\color{black}{ \displaystyle \log(0.00001)=\log(10^{-1})=-5}\) and on....

OpenStudy (solomonzelman):

and the closer the input approaches 0, the more magnitude the negative value gains

OpenStudy (anonymous):

ok thank you how do I get the range from it

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 1\times 10^{-\infty}}\) is a very very very small number (that apprahces zero and closer and closer to zero, the bigger this infity gets. right ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

that means, the closer x gets to zero, smaller value you are plugging in log(x)

OpenStudy (solomonzelman):

however when you plug this very small value into log(x) what what happens. \(\large\color{black}{ \displaystyle \log\left(x\right)}\) \(\large\color{black}{ \displaystyle \log\left( 10^{-\infty}\right)}\) \(\large\color{black}{ \displaystyle -\infty~\log\left( 10\right)}\) \(\large\color{black}{ \displaystyle -\infty}\)

OpenStudy (solomonzelman):

So, as far as negative values go, log(x) can get as 'small' as \(\large\color{black}{ \displaystyle -\infty)}\)

OpenStudy (anonymous):

so how would this effect my equation

OpenStudy (anonymous):

Isn't range just all the y values

OpenStudy (solomonzelman):

and can you get to the positive infinity range when having log(x) ?

OpenStudy (solomonzelman):

yes, range is all values of y

OpenStudy (solomonzelman):

I was trying to get first, to why this is so.

OpenStudy (anonymous):

yes just not 0 or negative right?

OpenStudy (solomonzelman):

it can be 0

OpenStudy (solomonzelman):

the range can be zero or negative

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \log(1)=0}\) (no matter what the base is) \(\large\color{black}{ \displaystyle \log(A)}\) when A<0 that gives a negative range

OpenStudy (anonymous):

ok but how do I use my equation to find the range